OpenStudy (goformit100):

Which of the following decimals has the LEAST value? a. 0.0013 b. 0.0103 c. 0.013 d. 0.13

OpenStudy (goformit100):

@elizabeth080808

OpenStudy (ajprincess):

\[0.0013=\frac{ 13 }{ 10000}\] \[0.0103=\frac{ 103 }{ 10000}\] \[0.013=\frac{ 130 }{ 10000}\] \[0.13=\frac{ 1300 }{ 10000}\] hope u can compare nw:)

OpenStudy (goformit100):

Ya. Thankyou Madam

OpenStudy (ajprincess):

welcome:)